The centrifugal air compressor of a gas turbine receives air



from the ambient atmosphere where the pressure is 1 bar and



the temperature is 300K. At the discharge of the compressor,



the pressure is 4 bar and the temperature is 480K and the



velocity is 100 m/s. The mass flow rate of air is 15kg/s. (a)



Determine the power required to drive the compressor and



(b) Explain why the given air exit temperature is more or less



than the reversible adiabatic compression temperature.

1
Expert's answer
2021-12-15T11:19:28-0500

Solution:



"Q\u2212W=\u0394h+\u0394K""\u0394h=(1)(480\u2212300)=180 \\frac{kJ}{kg}\n\u200b"

"\u0394K=0.5(100) ^2\n =5000 \\frac{J}{kg}\n\u200b\n =5 \\frac{kJ}{kg}"

"W=\u2212180\u22125=\u2212185 \\frac{kJ}{kg}\n\u200b""P=\u2212185(15)=\u22122775 kW"